OpenStudy (anonymous):

Identify the 12th term of a geometric sequence where a1 = 8 and a6 = –8,192.

OpenStudy (anonymous):

a6 = a1*r^5 -8192 = 8*r^5 r^5 = -1024 r = -2 so a12 = a1*r^11 = 8*(-2)^11 = -16384 i guess

OpenStudy (anonymous):

ahhhh, that not in my choice of answerss:(

OpenStudy (anonymous):

ahh I see now I got dizzy again lol it's too late here r = -4

OpenStudy (anonymous):

so a12 = 8*(-4)^11 sorry for many mistakes lol

OpenStudy (anonymous):

ahh, thats not good!:(

OpenStudy (anonymous):

but be sorry you have no idea how helpful you are(:

OpenStudy (anonymous):

yess thats right!

OpenStudy (anonymous):

what about Identify the 17th term of a geometric sequence where a1 = 16 and a5 = 150.06. Round the common ratio and 17th term to the nearest hundredth.

OpenStudy (anonymous):

a5 = a1*r^4 150.6 = 16*r^4 so u can figure r by using calculator 17th term to the nearest hundredth?? I don't understand my eng is not good enough lol

OpenStudy (anonymous):

haha ummm dont worry about that part? so like Identify the 17th term of a geometric sequence where a1 = 16 and a5 = 150.06. Round the common ratio and 17th term

OpenStudy (campbell_st):

stephanizer you need to evaluate the answer for your 1st term \[T _{12} = 8\times(-4)^{11}\] as the maths is correct.

OpenStudy (anonymous):

ok r^4 is approximately 6.4 so a17 = a1*r^16 = a1*((r^4))^4 = 8*(6.4)^4=.....

OpenStudy (anonymous):

but I want to understand eng too lol

OpenStudy (anonymous):

no no r is approximately 9.4 lol

OpenStudy (campbell_st):

I got \[150.6 = 16r^4\] r = 1.75 then \[T _{17} = 16 \times(1.75)^{16} =123802\]

OpenStudy (anonymous):

^^ how you figure r out ?? did you use calculator??

OpenStudy (campbell_st):

\[r^4 = \frac{150.6}{16}\] then \[r = \sqrt[4]{9.4125} = 1.7499\]